OpenStudy (anonymous):

The price of a sewing machine depreciates by 25/2% annually. If the machine was bought two years ago for Rs. 3600, what is its percent price ?

OpenStudy (anonymous):

find 52/2 % of 3600

OpenStudy (ganpat):

Present price = 3600 - 2 (52/2 % of 3600) as it is depreciating, it decreases and multiply by two, since it was bought 2 yrs ago..

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

it is not 52/2 but 25/2

OpenStudy (ganpat):

Present price = 3600 - 2 (25/2 % of 3600) so 25/2 % of 3600 = (25/2 ) * (1/100) * 3600 = 450 SO, equation changes to , 3600 - 2 (450)... calculate the answers, and re check my calculations..

OpenStudy (anonymous):

2700 ?

OpenStudy (ganpat):

perfect..... doing good. :)

OpenStudy (anonymous):

thanx.

OpenStudy (ganpat):

W/c..